OpenStudy (anonymous):

An unknown number b is 10 more than an unknown number k. The number b is also k less than 8. The equations to find k and b are shown below: b = k + 10 b = −k + 8 Which is a correct step to find k and b? Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis. Add the equations to eliminate k. Multiply the equations to eliminate b.

OpenStudy (texaschic101):

add the equations to eliminate k b = k + 10 b = -k + 8 ----------add 2b = 18 b = 9 now sub 9 in for b in either equation to find k b = k + 10 9 = k + 10 9 - 10 = k -1 = k check.. b = -k + 8 9 = - (-1) + 8 9 = 1 + 8 9 = 9 (correct)